My problem says to solve each equation in the real number system. The problem says 2x^3+3x^2+2x+3=0. I factored the problem by grouping and came out with (2x^2+3)(x^2+1)=0, but when I solve for x I come out with imaginary numbers and I am not supposed to. I would apprecaite any help you could give me in solving this problem. Thanks, Britta
Answer by venugopalramana(3286)

My problem says to solve each equation in the real number system. The problem says 2x^3+3x^2+2x+3=0. I factored the problem by grouping and came out with (2x^2+3)(x^2+1)=0,
NO...DID YOU TYPE WRONGLY?2X+3 IS A FACTOR NOT 2X^2+3
X^2(2X+3)+1(2X+3)=(2X+3)(X^2+1)=0...SO THERE IS ONE SOLUTION IN REAL NUMBERS NAMELU...2X+3=0...OR...2X=-3...OR...X=-3/2=-1.5
but when I solve for x I come out with imaginary numbers and I am not supposed to. I would apprecaite any help you could give me in solving this problem. Thanks, Britta